A body tied to a weightless thread is completely immersed in liquid. For a liquid of density 0, 6 g/cm3 the thread tension is equal to 2N while for a liquid of density 0, 9 g/cm3 the thread tension is equal to 1, 7 N. Determine the body density. The answer should be expressed in [g/cm3].
1
Expert's answer
2020-11-25T10:51:41-0500
"T_1=\rho gV-\rho_1 gV\\T_2=\rho gV-\rho_2 gV\\\frac{T_1}{T_2}=\frac{\rho -\rho_1}{\rho -\rho_2 }\\\frac{2}{1.7}=\frac{\rho -0.6}{\rho -0.9 }\\\rho=2.6\frac{g}{cm^3}"